### v1

# A Video is Worth 10,000 Words: Training and Benchmarking with Diverse Captions for Better Long Video Retrieval

Abstract
Existing long video retrieval systems are trained and tested in the paragraph-to-video retrieval regime, where every long video is described by a single long paragraph. This neglects the richness and variety of possible valid descriptions of a video, which could be described in moment-by-moment detail, or in a single phrase summary, or anything in between. To provide a more thorough evaluation of the capabilities of long video retrieval systems, we propose a pipeline that leverages state-of-the-art large language models to carefully generate a diverse set of synthetic captions for long videos. We validate this pipeline’s fidelity via rigorous human inspection. We then benchmark a representative set of video language models on these synthetic captions using a few long video datasets, showing that they struggle with the transformed data, especially the shortest captions. We also propose a lightweight fine-tuning method, where we use a contrastive loss to learn a hierarchical embedding loss based on the differing levels of information among the various captions. Our method improves performance both on the downstream paragraph-to-video retrieval task (+1.1% R@1 on ActivityNet), as well as for the various long video retrieval metrics we compute using our synthetic data (+3.6% R@1 for short descriptions on ActivityNet).
**footnotetext: Work performed during internship with SRI International.
1 Introduction
Figure 1: In real-world text-to-video retrieval, users could use diverse queries. Standard long video datasets use only paragraph-style captions (“Existing”, “Full paragraph”), which does not allow for training or evaluation on a representative set of long video descriptions. Practical applications also require the ability to handle complex, short, and partial descriptions of a long video. In this work, we introduce an approach to generate, evaluate, and train on such diverse video description data.
If a picture is worth 1,000 words, then a video is worth 10,000. Consider the variety of possible captions for a video, illustrated in Figure 1, which describe the video at varying levels of detail. These can vary substantially in terms of semantics and structure. Nevertheless, a video retrieval system ought to be able to map any of these captions to the video, whether they are short, long, abstract, or even describe only part of the video. We thus seek to redefine the video retrieval problem, where a system, given some text, must retrieve a video.
Table 1: 10k Data and Notation. We give our diversity axes, levels, and show an example of the captions for 1 video.
Axis Level Example
Ground Truth Full (f) People are sitting in kayaks paddling in the water. They go under a rock and through a tunnel.
Partial (p) They go under a rock and through a tunnel.
Summarization Short (s) Kayakers paddle, go under rock, through tunnel.
Medium (m) People in kayaks paddle, pass under rock, navigate through tunnel in water.
Long (l) A group of kayakers paddle through water, passing under a rock and navigating through a tunnel.
Simplification Elementary (e) People are in small boats and paddle in the water. They go under a big rock and through a tunnel.
Intermediate (i) Individuals are seated in kayaks, using paddles to navigate through the water. They pass beneath
a large rock formation and venture through a tunnel.
University (u) A group of individuals are situated in kayaks, propelling themselves forward with paddles as they traverse
the water. They maneuver beneath a substantial rock structure and proceed through a tunnel.
Video retrieval has long been a important interdisciplinary task that connects the state-of-the-art in natural language processing and computer vision [25, 34]. At its core, solving the problem requires a system that understands both video and text. However, current video retrieval literature often considers just short clips, which cannot be described by such a variety of captions, and thus obscures the problem. Increasingly more work does focus on long videos, but it uses only full paragraphs for retrieval [7, 6], neglecting to consider the rich space of valid video captions. However, the real world includes long videos with multiple events, and these admit many vague, abstract, or partial descriptions a person (e.g., doing video search) might give. In practice, the strict mapping of a single caption to a single video in instance-based video retrieval neglects both the potential ambiguities of the caption as well as the fact it is valid to describe the same video in a variety of ways [48].
To address this we formulate the 10k Words problem, a novel video retrieval setting which includes diverse descriptions generated for long videos with multiple events. This problem identifies key axes of variation, including simplification, summarization, and duration, that we use to curate pools of captions with non-trivial differences in structure and semantics. The problem introduces challenging ambiguities, since some captions will not mention all the details that distinguish a video from similar, related videos.
We instantiate the 10k Words problem by building 3 datasets, ActivityNet10k, QuerYD10k, and LF-VILA10k, by augmenting existing datasets [17, 37, 45] with diverse captions. These augmentations are only possible given the flexibility and accuracy of recent large language models (LLMs) [8], which we combine with some simple automatic manipulations to synthesize the diverse 10k Words datasets. We provide a thorough definition of the 10k Words problem and our data generation pipeline in Section 2.
These datasets can help us detect failures of existing models to capture the space of text descriptions as well as help us to mitigate those failures, and we show both. For detection, we consider a representative set of state-of-the-art video models and show that they struggle to adequately solve the 10k Words problem in Section 3, struggling especially with short, summary-style captions. We then demonstrate the effectiveness of a simple mitigation strategy that uses 10k datasets to augment standard datasets during training. This can provide an inexpensive boost to performance on both the 10k datasets and the original standard datasets, or be used to increase data efficiency. However, simple data augmentation does not fully leverage the knowledge that multiple possible captions correspond to the same video despite meaningful structural and semantic differences.
We thus propose an approach that uses 10k captions to guide the text embedding space. Prior work has pointed out that some paragraph captions may already be valid descriptions of multiple videos [48], and this is only more frequent with our synthetic short summary captions. To help our model address this we finetune it to not only bring synthetic caption features and real video features close to each other, but also to align a projection of the synthetic text features with real text and real video features. The projection step allows the model to learn embeddings that account for the different levels of information between the synthetic caption, the real caption, and the real video. We achieve SOTA performance on 10k Words while also boosting performance on the standard paragraph-to-video retrieval task. In summary, we contribute the following:
  • We formulate and propose the 10k Words problem, a framework for characterizing the broad spectrum of valid descriptions for long videos.
  • We create ActivityNet10k, QuerYD10k, and LF-VILA10k by using our flexible data generation pipeline.
  • We benchmark SOTA models in a zero-shot fashion, and reveal that they struggle on the 10k Words problem.
  • We leverage 10k data and lightweight contrastive losses to achieve SOTA performance for 10k Words, with a +1.1% R@1 improvement on standard ActivityNet retrieval and +2.8% R@) improvement on ActivityNet10k.
2 10k Words Dataset
2.1 10k Definition and Generation
Table 2: Automatic dataset statistics. We show the average change in unique nouns and verbs, as well as word count and length for the different dimensions of our 10k Words ActivityNet datasets.
Summarization Simplification Summarization and Simplification
Metric
Source Short Medium Full Length Primary Secondary University S and P S and S S and U
$\Delta$ Nouns
-5.37 -1.69 0.06 -1.16 0.97 3.84 -6.20 -5.83 -5.41
$\Delta$ Verbs
-5.01 -1.97 -0.77 -0.65 0.90 1.95 -5.19 -5.02 -4.84
Word Count
49.77 8.77 29.29 37.39 43.54 48.31 56.13 8.50 9.28 10.75
Word Length
5.09 6.10 5.40 5.51 4.97 5.49 5.97 5.27 5.75 6.10
We now describe the procedure for creating a dataset of videos and natural language description queries such that the queries cover many of the possible ways of describing a video. We start with an existing dataset of annotated videos and enrich the set of descriptions in the dataset, so it covers more of the possible ways to describe the videos. To cover the broadest possible spectrum of natural language queries for a video we start by defining three axes along which a video’s description can vary: duration, summarization, and simplification. The first dimension, duration, refers to how many of the events in a video are described by a given query. The last two dimensions cover different ways of using language to describe the same video. Prior to LLMs these last two dimension would have been difficult to simulate effectively and reliably, and perhaps would have required expensive human annotations. However, we are able to effectively prompt LLM to gather such data [8, 57]. For examples of these dimensions, see Table 1.
More precisely, consider a long video with $E$ events ${e_{1},e_{2},\ldots,e_{E}}$. Each event $e_{i}$ has a corresponding short video clip $v_{i}$ and natural language description of that clip $t_{i}$ with the set of clips and texts being denoted $V$ and $T$. As such, the original long caption could be a paragraph, long sentence, or, more typically, the concatenation of video segment captions. This is typical for the datasets shown in Table 3. Each axis of the 10k problem implements a function that takes such a video with event segmentation and descriptions as input and outputs a new augmented version of the same video with a new set of segments and descriptions. We now cover each of these functions.
Summarization. Descriptions of videos can vary in length, where at one extreme they describe every detail in the video, at the other they briefly describe the main idea leaving out some significant details, and in between relevant details are progressively grouped and redundant elements are pruned. At one end of this spectrum a video retrieval model must be able to parse details and at the other end it must be able to understand a gestalt. To augment a video on this axis we prompt an LLM with the ground truth descriptions $T$ (concatenated) and instruct it to generate summaries. If the concatenated description has $L$ words then we ask the LLM to generate three summaries with $\lfloor{L}\cdot\frac{l}{7}\rfloor$ words each for $l\in\{1,4,7\}$. At full length ($l=7$) this should just re-phrase the concatenated caption, but at smaller lengths the LLM must leave out information. We observe that GPT-3.5***gpt-3.5-turbo-0613 is able to achieve close to the desired word count most of the time. This only changes $T$, leaving $E$ and $V$ unchanged. We share exact prompts in the appendix.
Table 3: Base Datasets which we use for our synthetic caption generation pipeline. We use only the ‘train’ and ‘val-1’ splits of ActivityNet, and do some additional filtering for extremely long outlier captions from LF-VILA and QuerYD.
Dataset # Videos Video Len (s) Text Len (w) Duration (h)
ActivityNet [27] 14926 117.5 49.8 487
LF-VILA-7k [54, 45] 7020 203.9 155.4 397
QuerYD [37] 2474 264.3 203.8 181
Simplification. Descriptions of videos can vary in terms of their conceptual simplicity, where an idea could be described at the level of a college graduate, or else simplified for a kindergartener, and a good retrieval model should map all these descriptions to the same video. We capture this dimension by providing an LLM with the same ground truth description as for summarization and instruct it to output a simplified version. This is done for three levels of reading comprehension described to the LLM as “primary school”, “secondary school”, or “university” reading level. This augmentation only modifies $T$, leaving $E$ and $V$ unchanged. We share exact prompts in the appendix.
Duration. Descriptions of videos can be partial, intending to cover only a segment of the video, but the video should still be retrieved when these are used as queries. In our dataset we implement this by choosing a contiguous subset of events $\tilde{E}={e_{i},\ldots,e_{j}}$ with start and end index $i$ and $j$. The corresponding set of video clips $\tilde{V}$ and captions $\tilde{T}$ are selected to create the augmented video.
Table 4: Meaning Preservation results. For each item, we present the annotators with a paragraph and three synthetic captions: one generated from the paragraph, one from a neighbor, and one at random. We show how often each caption is judged as a match to the paragraph.
Table 5: Simplification Validation results. For each paragraph we ask the annotators to rank the three synthetic simplification captions from simplest to most complex. The majority results show that actual complexities correlate well with the intended simplification.
Table 6: Hallucination Prevalence results. We treat each potentially hallucinated word in the generated caption as an item, and show results over all votes, as well as the majority label for each item. This suggests most potential hallucinations are actually consistent with the source caption.
Different Unsure Matches
Actual Match 0% 4% 96%
Neighbor 20% 28% 52%
Random 100% 0% 0%
Simplest Middle Most Complex
Elementary 84% 16% 0%
Intermediate 16% 84% 0%
University 0% 0% 100%
Different Unsure Matches
Total 24.75% 8.08% 67.17%
Majority (per-word) 18.18% 9.09% 72.73%
Table 5: Simplification Validation results. For each paragraph we ask the annotators to rank the three synthetic simplification captions from simplest to most complex. The majority results show that actual complexities correlate well with the intended simplification.
Table 6: Hallucination Prevalence results. We treat each potentially hallucinated word in the generated caption as an item, and show results over all votes, as well as the majority label for each item. This suggests most potential hallucinations are actually consistent with the source caption.
Table 7: Unanimous annotator agreement, or the portion of items per section for which all annotators give the same label.
Meaning Preservation Simplification Validation Hallucination Prevalence
Random 11.11% 11.11% 11.11%
Actual 71.67% 64.00% 55.00%
10k Datasets. We combine these axes to construct 10k versions of ActivityNet, QuerYD, and LF-VILA-7k (Table 3). We construct our 10k Words datasets by taking the per-segment captions available for the datasets described in Table 3 and feed them to GPT-3.5 with relevant prompts. Starting from each video in a base dataset like ActivityNet, we include 11 captions for each video: 1 full caption (original ground truth paragraph), 3 captions for the levels of simplification (elementary, intermediate, and university), 3 captions for the levels of summarization (short, medium, and long), 3 captions that combine summarization and simplification by generating simplifications for the short summaries, and 1 caption corresponding to a random subset of the original video segments by duration augmentation. We show examples for summarization, simplification, and partial captions, and introduce relevant shorthand, in Table 1.
We refer to the 10k Words version of LF-VILA-7k, a sample from the original LF-VILA [45], as LF-VILA10k. We use all of QuerYD as a validation set, since the initial small size of its validation set makes it challenging to distill useful insights, and create QuerYD10k. We create ActivityNet10k for ActivityNet. We provide details on LLM prompts and costs in the appendix.
2.2 Dataset Analysis
We provide some fine-grained statistical measures to examine the nature of our generated data in Section 2.2.1. We also perform a study on a sample of our data using human annotators to further validate the claims regarding our data and ensure that it is free from undesirable artifacts (e.g. hallucinations) in Section 2.2.2.
2.2.1 Automatic Analysis
First, we perform an automatic analysis of our generated data. For the sake of brevity we focus on ActivityNet, with metrics for other datasets provided in the supplement. From Table 2, note that summarization and elementary level simplification tend to remove nouns and verbs ***Extracted using https://spacy.io part-of-speech tagging, while intermediate and university levels tend to add nouns and verbs. Also note the word counts, where summarization, as expected, reduces the average number of words, while simplification to elementary level reduces the average length of words.
Table 8: Text to Video retrieval performance for our benchmark on ActivityNet. First we reproduce results for standard paragraph to video retrieval. Then, we give the average performance on short 10k Words captions, long 10k Words captions, and partial captions. We use the standard recall at top-1 metric (R@1) as well as the average of recall at top-1/5/10 (Avg. Recall). We explain how we aggregate 10k Words captions as Full, Short, Long, and Partial in Section 3.1. We explain the finetuning methods in Section 4.
Model Standard 10k Words
ANet Full QuerYD Full LF-VILA Full ANet10k All ANet10k Short ANet10k Long ANet10k
R@1 Avg. R R@1 Avg. R R@1 Avg. R R@1 Avg. R R@1 Avg. R R@1 Avg. R R@1 Avg. R
zero-shot VideoCLIP 6.3 16.2 7.4 15.9 5.1 11.1 5.3 13.9 4.0 10.9 6.2 16.3 6.6 16.3
Frozen 14.0 32.4 13.7 27.5 26.1 43.4 11.4 27.4 8.7 22.8 14.2 32.0 11.0 27.3
COSA 34.3 55.7 34.4 49.6 66.8 78.4 23.9 43.4 16.2 33.9 31.7 52.9 23.7 43.4
InternVideo 47.9 68.2 50.1 63.6 49.5 63.5 36.0 57.5 27.8 49.3 44.4 65.9 35.0 56.8
finetune Domain 58.8 77.7 - - 95.9 98.6 42.0 62.8 29.9 51.4 53.8 73.8 43.2 64.5
10k words 59.3 78.1 - - 97.4 99.1 44.3 65.2 32.3 54.8 56.2 75.5 44.3 65.9
Ours 59.9 78.5 - - 97.5 99.1 44.8 66.1 33.5 56.3 56.2 75.9 44.7 66.2
2.2.2 Annotator Analysis
To validate the fidelity and utility of our captions we recruit 15 human annotators to examine our captions in an IRB approved study. We design a survey that consists of 3 sections, according to the properties of the data we wish to examine. In the first section, we analyze whether the LLM-generated captions preserve original meaning. In the second section, we ensure that when the LLM performs the simplification in a manner that is meaningful to humans. In the third section, we verify the extent to which hallucinations occur in the LLM-generated captions. Each section has 5 questions. We divide our annotators into groups of 3, to allow for analysis of inter-annotator agreement, and thus distribute 5 versions of the survey, covering a sample of 75 videos from the validation set of ActivityNet. Next, we provide more detail regarding the design of the survey and results for each section. As evidence of the survey’s validity, we show inter-annotator agreement in Table 7. For full details, including exact surveys, please see the appendix.
Meaning Preservation. For each question in this section, we randomly sample one real caption, and assign it to be the “ground truth” caption. We then sample 3 generated captions – one generated from the “ground truth” caption, one generated from the “ground truth” caption’s nearest neighbor caption***Maximum cosine similarity between “ground truth” captions using OpenAI text-embedding-ada-002 embedding., and one generated from a random unrelated caption. We ask the annotator to determine, for each of the 3 generated captions, whether they believe it describes the same video as the ground truth caption. We observe, in Table 6, that the generated caption for any given source caption is very consistently judged to be from the same video, with neighbors being determined to come from the same video to a lesser extent.
Simplification Validation. For each question in this section, we randomly sample one real caption and show the annotator the “primary school”, “secondary school”, and “university” captions generated by the LLM. Then we ask the annotator to rank them from most to least complex. In Table 6, we find very little ambiguity in the simplification rankings. Annotators consistently judge “university” to be the most complex, and “primary school” to be the least complex.
Hallucination Prevalence. For each question in this section, we sample some real caption and one of its generated captions, either the full length summary, or one of the full length simplification captions. We then use spaCy***https://spacy.io part-of-speech tagging to extract the nouns and verbs which appear in the generated caption but not the original. We then present both captions to the annotator, and for up to 3 of the potentially hallucinated words, we ask whether or not they change the meaning of the original caption. We find, in Table 6, that of the new words for the generated captions, annotators tend to judge that they typically correspond to entities and actions that are already depicted in the source captions. This, along with the results from the first section of the study, suggest that the prevalence and impact of potential hallucination is quite limited.
All this automatic and human analysis leads us to conclude that the captions we generate in our novel paradigm are diverse, robust, and, as we will continue to demonstrate, very useful.
3 Benchmark and Analysis
Figure 2: We compute the difference in R@1 between retrieval with full paragraph captions, compared to short (s), medium (m), long(l), elementary (e), intermediate(i), university (u), and partial captions and show the difference as a percentage relative to the full paragraph R@1. Each bar is the average of these percentages for all 3 datasets. Notice that while deltas are consistent between medium and long captions, there is a drop off for short captions and partial captions, demonstrating that a key facet of this problems difficulty is the introduction of ambiguity due to missing information.
Figure 3: We measure the retrieval for the datasets across all zero-shot models and captions. We plot the number of videos which are retrieved as the top-1 result for the given number of model-caption pairs. At $x=0$, this is the number of videos which are not retrieved by many model, for any caption. At $x=40$, this is how many videos are retrieved as the first result for all 4 models, for all 10 possible captions.
3.1 Preliminaries
Models
For our 10k Words benchmark, we evaluate the performance of VideoCLIP [50], Frozen [7], InternVideo [46], and COSA [12] as a set of representative video-language models. For COSA we use the ‘itm’ retrieval and for InternVideo we use the dual softmax. We evaluate these in a zero-shot manner to show 10k data is difficult for existing models, and show that both finetuning and our proposed approach can improve performance in Section 4.
Experiment Details
We run our experiments on nodes containing between 1 and 4 GTX 2080Ti, RTX A5000, and RTX A6000 GPUs, depending on the demands of each model. When training, we follow the settings provided in the publicly available code of the models we chose. Please see our code (with our dataloaders and other relevant modifications) for reproducibility.
Table 9: We provide further zero shot results (R@1) for LF-VILA10k and QuerYD10k.
QuerYD10k LF-VILA10k
Model All Short Long Partial All Short Long Partial
VideoCLIP 6.8 7.0 6.4 7.8 4.3 4.1 4.4 5.0
Frozen 13.3 12.2 15.1 10.5 21.4 17.7 25.6 19.3
COSA 27.8 27.4 29.6 21.9 42.8 41.0 44.0 45.4
InternVideo 46.8 44.8 49.2 45.3 39.9 35.4 45.2 37.1
Metrics
We primarily use text-to-video recall @ K (R@K) to measure performance. Given a list of text queries and video targets relative to a database of videos to be retrieved, R@K measures the percentage of queries for which the ground truth target was retrieved at rank K. We use this to measure overall performance, but also to measure subset performance on Summarization and Simplification by considering the appropriate subsets of captions.
To measure performance on the Duration axis we consider whether the partial and full captions can retrieve the full length video. The “Full” setting measures how often the full caption (f) retrieves the video at rank K or better, which represents performance as measured by the standard datasets. Since we use the standard ActivityNet settings, these can be compared with numbers from other papers; however, since we use unique splits for QuerYD and LF-VILA, these numbers are not comparable. The “Partial” setting measures how often the partial caption (p) retrieves the same full length video. The “Short” setting measures performance of full length video retrieved by short captions including short summarization (s) and simplifications of it (s+e, s+i, s+u). Similarly, we also report performance on the “Long” setting which include long summarization (l) and simplifications of it (l+e, l+i, l+u). The “All” setting is an average of the Partial, Short, and Long, weighted corresponding to the number of captions that comprise each.
3.2 Zero-Shot Results
We provide the zero-shot benchmarking results in Table 8, with remaining results on the LF-VILA and QuerYD in Table 9. The methods that perform best for Full paragraphs also tend to perform best for the Long, Short, and Partial captions. Notably, there is only a minor gap in retrieval performance between the Full and Long captions, with a larger difference between Full and Partial captions, and a significant drop for the Short captions. We also show using Figure 2 that the shortest summaries are the hardest to retrieve. Note that the differences in Figure 2 are computed as a percentage of the Full R@1. With this in mind, we see that COSA seems to be by far the least robust to 10k Words data, with the largest relative changes in performance. By contrast, VideoCLIP and especially Frozen are often benefited by the 10k Words data, particularly when the axis is rewording the caption (e, i, u) rather than removing it (s, p).
With these interesting performance discrepancies in mind, we leverage our suite of models and 10k Words data as a way to evaluate the retrieval difficulty of each dataset. To do so, we consider all 10 possible captions (we exclude Partial) and a model as 10 model-caption pairs, so with 4 models, we have 40 pairs. We then measure, for each video in a dataset, how often (for the 40 pairs) it is retrieved as the first-ranked result. We plot this in Figure 3 and find that LF-VILA tends to be the easiest dataset, with QuerYD having a significant portion of videos that are not retrieved correctly no matter which model and caption are used.
4 Long Video Finetuning
Figure 4: We perform contrastive finetuning for retrieval with video-caption pairs.  One key novelty is the project we propose for 2 projection losses, between projected 10k  text features and both ground truth text features as well as video features, respectively.  The other key novelty is the data augmentation, where we compute standard contrastive loss but each caption is sampled  randomly from the complexity distribution for that entity, according to a mixing ratio, $\eta$.
In this section we propose a method for finetuning pre-trained models to retrieve videos from any relevant caption. We then highlight results.
4.1 Method
We propose losses complementary to any model that uses contrastive-style loss, in our case, finetuning both COSA and InternVideo. We illustrate this approach in Figure 4. We sample a batch of videos with corresponding captions and apply a loss that pushes matching video and caption embeddings closer together. To encourage the model to associate all of the descriptions for a video with that video we add the additional 9 captions for a video during training.
Figure 5: We measure the extent to which our data augmentation improves performance in the data constrained regime, training only with the indicated amounts of data, and computing paragraph-to-video and 10k Words retrieval for the resulting trained models. We show, for ActivityNet10k, that our finetuning method (ours) is consistently better than vanilla COSA finetuning (C) for retrieval with Full, Long, and Short captions.
[…]
$L=l_{\text{itc}}+\alpha_{\text{t2t}}\cdot l_{\text{t2t}}+\alpha_{\text{proj}}\cdot l_{\text{proj}}$ (1)
We use 3 settings in this paper: “Domain Finetune” which is just the default setting of whichever model (COSA or InternVideo) we are finetuning, “10k Finetune” where we set $\eta=0.75$ but keep $\alpha_{\text{t2t}},\alpha_{\text{proj}}=0$, and then “Ours” where we set $\eta=0.75$ and $\alpha_{\text{t2t}},\alpha_{\text{proj}}=0.1$. We set these values after some ablations, which we report in the supplement for reference.
Table 10: LFVILA COSA finetuning results.  See Section 4.1 for explanations of settings. 10k Words data helps performance.
Finetune Method All Short Long Partial
R@1 Avg. R R@1 Avg. R R@1 Avg. R R@1 Avg. R
Domain 77.3 86.9 65.2 78.4 90.2 95.9 73.8 84.9
10k 85.2 92.6 78.2 89.2 95.3 98.2 73.0 83.9
Ours 84.9 92.5 78.0 89.1 95.0 98.1 71.9 83.3
Table 11: ActivityNet InternVideo finetuning results. Our method improves performance here as well, although the majority of the improvement here can be attributed to the data.
Finetune Method All Partial Short Long
R@1 Avg. R R@1 Avg. R R@1 Avg. R R@1 Avg. R
Domain 41.4 62.7 52.1 72.4 29.1 51.2 48.5 69.7
10k 42.2 63.6 52.6 72.8 30.0 52.6 49.1 70.5
Ours 42.2 63.7 52.6 72.9 30.0 52.7 49.4 70.7
4.2 Results
Table 8 shows the results for COSA finetuning on ActivityNet (”Domain”). We observe that finetuning by sampling from 10k Words data (”10 Words”) yields considerable improvements for retrieving with 10k Words captions. We also see that introducing our novel losses (”Ours”) further improves performance. We show that these findings hold when finetuning on other datasets, such as LF-VILA (Table 10), as well as when we adapt our data sampling and losses for another state-of-the-art model, InternVideo, in Table 11. This also reveals that the LF-VILA dataset is easier to tackle once the model can finetune. This is perhaps an artifact of LF-VILA’s original automatically generated captions, or due to the fact that LF-VILA is open domain and lacks the fine-grained differences present in ActivityNet.
Table 12: Data sampling ablation. We experiment with using 10k Words data consisting of the indicated types: Simplification (l+e, l+i, l+u), Joint (s+e, s+i, s+u), and/or Partial. We observe that even when the model is trained only with the simplification captions, it has generalizes well for types of captions.
Sampled 10k Types ActivityNet
Simplification Joint partial Full All Short Long
59.3 43.55 30.7 56.4
60.1 45.1 33.2 57.0
58.1 43.1 30.8 55.3
58.6 44.3 33.4 55.1
57.7 43.5 32.7 54.3
58.5 44.2 32.8 55.5
Since part of our contribution is a data augmentation strategy, we also evaluate its performance by finetuning with different fractions of the original ActivityNet data in Figure 5. We find that using 10k Words data consistently improves performance for all amounts of training data.
4.3 Ablations
In Table 12, we find that the finetuning is quite sensitive to the type of data sampled, where using partial data while training our losses tends to be detrimental to the performance, regardless of which caption is used. Notice how finetuning even with only simplification samples, which are paragraph length, results in +0.8% improvement for Short R@1 compared to “Domain” finetuning, which suggests the 10k Words data has nice properties for generalization. We perform ablations for $\eta$, $\alpha_{\text{proj}}$, and $\alpha_{\text{t2t}}$ as well (see appendix, and find that most hyperparameter changes are not more significant than variance across runs, except for setting $\eta=0$, which is detrimental to 10k Words performance.
5 Related Work
Video-Language Models. Recent efforts have started to focus on building a multimodal foundation model that understands not just vision and language, but captures the extra complexities of video and language. These approaches are typically pre-trained in a manner inspired by CLIP [39] and ALIGN [24], using sets of video and text pairs with varying levels of noise [34, 7, 60], where the text is some form of video description, caption or summary. These models are typically pretrained on some task or set of tasks, and then used on downstream tasks either in a zero-shot manner or after finetuning. Early approaches use pre-computed features to represent videos [59]. Typical models learn a shared embedding space between the videos and text [50, 4, 16, 18, 34, 35, 23]. Others process concatenated video and text inputs with cross-modal encoders [44, 30, 43, 14, 15, 62, 49]. Some even use still images or average frame embeddings and achieve quite strong performance [9, 29, 6]. However, as computational resources have scaled, so have the methods, and many current approaches learn to compute features from raw video [7, 46, 12, 31]. Along with this trend models are branching out from contrastive learning to incorporate other learning tasks as well, even generative objectives including captioning [46, 12, 51, 13, 21, 11, 55, 28]. In this paper we focus on the video retrieval task, and show results using VideoCLIP [50], Frozen [7], COSA [12], and InternVideo [46] as a representative set of video-lanuage models.
Long Video Understanding. Videos in the computer vision literature tend to be short – the average length of videos in tentpole datasets [41, 52, 19, 10, 22] is under 30 seconds. Over the years, some have introduced datasets consisting of longer videos [17, 5, 61, 37, 45]. With the introduction of these datasets, larger GPUs, and advancements in vision-text modeling, many researchers have begun proposing methods that either address long video as a first-class interest [6, 45, 40], or at least, are flexible for both long and short videos [31, 11, 13, 46, 12]. In this paper we focus on long video from the perspective of the retrieval task, and we propose a method for data expansion to allow for more thorough development, evaluation, and benchmarking of long video understanding approaches. This is reminiscent of [47], however, unlike their work, we generate novel synthetic captions, and restrict our training and analysis to the more common single query retrieval setting.
Text Summarization. Our work bears some resemblance to efforts in the areas of controllable text summarization and simplification. For these tasks, control tokens dictate how a model simplifies or summarizes text while preserving its meaning. The definition of these tokens is a key differentiator between papers [2] – they can be user-defined [56, 36, 3, 1, 58], or optimized over some data [33, 32, 38, 26, 42]. Additionally, these pipelines often involve a human-in-the-loop at inference time to give keywords [20] and can require heavy labeling [53, 2]. We opt to use our LLM approach to avoid a reliance on control token definition, human-in-the-loop inference, or depedency on strongly annotated data and the biases that come from training on such limited data. As a result, our outputs are somewhat less “controllable” but are, as shown in Section 2, sufficient for our needs.
6 Conclusion
We showed how the video retrieval problem can be expanded to capture all ways someone might describe a video. We built ActivityNet10k, QuerYD10k, and LF-VILA10k, whose captions refer to videos using a variety of natural language descriptions instead of just one. We showed that SOTA models fail to generalize well to all potentially valid descriptions, and proposed a novel finetuning approach that can retrieve videos effectively from a range of descriptions. We proposed a finetuning approach to solve that more fine-grained retrieval problem more accurately. This work shows how the complexities of video understanding can be measured and addressed using language as a guide. We hope future work will continue to explore the complete spectrum of language that can be used to capture video content.
Supplementary Material
7 GPT-3.5 Details
7.1 Prompts and Costs
We share prompts for summarization, simplification, and the combination of the two (joint). In the main paper, summarization is denoted as s, m, l depending on length, where s has 1 word and m has 4 words for every 7 words in l. Simplification is denoted by l+e, l+i, l+u. Joint is s+e, s+i, s+u.
We reduce the cost in terms of input token counts by batching our inputs. For example, we are generating 3 different summarizations per paragraph, but the source paragraph is the same in all 3 cases. So, instead of passing the input once for each level of summarization (3 times total), we pass the input once, and ask for all summarizations to be present in the output, reducing our input tokens by a factor of 3. So, our final prompts are as follows for summarization, simplification, and joint:
Summarization
You are a helpful writing assistant, with a speciality in summarizing text-based scene descriptions. You will be asked to write 3 summaries of the scene described in the following paragraph, indicated by PARAGRAPH. Do not modify the indicated order of events. Prioritize visual details. Do not hallucinate. Do not describe objects or events that do not appear in the original paragraph. PARAGRAPH: ORIGINAL PARAGRAPH. Label this summary as SUMMARY_1. For this summary, please write 10 words which summarize the scene described by the PARAGRAPH. Do not use more or less than 10 words. Without using more than 10 words, write complete sentences. Label this summary as SUMMARY_4. For this summary, please write 40 words which summarize the scene described by the PARAGRAPH. Do not use more or less than 40 words. Without using more than 40 words, write complete sentences. Label this summary as SUMMARY_7. For this summary, please write 70 words which summarize the scene described by the PARAGRAPH. Do not use more or less than 70 words. Without using more than 70 words, write complete sentences.
Simplification
You are a helpful writing assistant, with a speciality in simplifying and rewriting descriptions for different age groups and reading levels. You will be asked to write 3 versions of the scene described in the following paragraph, indicated by PARAGRAPH. Do not modify the indicated order of events. Prioritize visual details. Do not hallucinate. Do not describe objects or events that do not appear in the original paragraph. PARAGRAPH: ORIGINAL PARAGRAPH. Label this version as VERSION_primary_school. For this version, rewrite the PARAGRAPH with 70 words to make it suitable for a primary school reading level. Label this version as VERSION_secondary_school. For this version, rewrite the PARAGRAPH with 70 words to make it suitable for a secondary school reading level. Label this version as VERSION_university. For this version, rewrite the PARAGRAPH with 70 words to make it suitable for a university reading level.
Joint
You are a helpful writing assistant, with a speciality in summarizing text-based scene descriptions. You also have a speciality in simplifying and rewriting descriptions for different age groups and reading levels. You will be asked to use 10 words to write 3 summaries of the scene described in the following paragraph, indicated by PARAGRAPH. Do not modify the indicated order of events. Prioritize visual details. Do not hallucinate. Do not describe objects or events that do not appear in the original paragraph. PARAGRAPH: ORIGINAL PARAGRAPH. Label this version as VERSION_primary_school. For this version, rewrite the PARAGRAPH with 10 words to make it suitable for a primary school reading level. Do not use more or less than 10 words. Without using more than 10 words, write complete sentences. Label this version as VERSION_secondary_school. For this version, rewrite the PARAGRAPH with 10 words to make it suitable for a secondary school reading level. Do not use more or less than 10 words. Without using more than 10 words, write complete sentences. Label this version as VERSION_university. For this version, rewrite the PARAGRAPH with 10 words to make it suitable for a university reading level. Do not use more or less than 10 words. Without using more than 10 words, write complete sentences.
7.2 Automatic Analysis
Table 13: Automatic dataset statistics for LF-VILA10k. We show the average change in unique nouns and verbs, as well as word count and length.
Summarization Simplification Summarization and Simplification
Metric
Source Short Medium Full Length Primary Secondary University S and P S and S S and U
$\Delta$ Nouns
-11.77 -4.23 1.49 -1.40 3.75 11.14 -14.35 -13.18 -12.36
$\Delta$ Verbs
-2.60 0.90 4.32 1.96 7.63 11.71 -3.07 -2.34 -1.86
Word Count
155.40 36.06 76.30 105.43 129.52 136.58 154.13 28.94 31.85 34.83
Word Length
4.66 5.00 4.96 5.18 4.79 5.25 5.74 4.66 4.90 5.12
Table 14: Automatic dataset statistics for QuerYD10k. We show the average change in unique nouns and verbs, as well as word count and length.
Summarization Simplification Summarization and Simplification
Metric
Source Short Medium Full Length Primary Secondary University S and P S and S S and U
$\Delta$ Nouns
-27.25 -20.38 -14.81 -14.83 -9.26 -2.55 -32.21 -31.16 -29.27
$\Delta$ Verbs
-12.10 -7.90 -4.46 -3.04 1.26 4.56 -14.11 -13.57 -12.83
Word Count
207.86 53.41 86.69 114.55 150.97 164.26 181.92 34.81 37.28 43.10
Word Length
5.47 5.89 5.72 5.79 5.27 5.66 6.02 5.37 5.73 5.98
We provide LF-VILA and QuerYD to complement Table 2 in Table 13 and Table 14, respectively. These are conistent with the major trends for ActivityNet10k, with the notable difference that since these captions are longer, the absolute differences are larger.
7.3 Annotator Analysis
For an example survey, please refer to the attached material.
8 Ablations
Table 15: Projection ratio and text to text ratio ablations.
Hyperparameter ActivityNet
$\eta$ $\alpha_{\text{proj}}$ $\alpha_{\text{t2t}}$ Full Short Long
0.75 0.1 0.1 59.9 33.5 56.2
0.75 0.1 0.0 59.3 33.2 56.5
0.75 0.1 0.2 59.6 33.3 56.5
0.75 0.1 0.5 59.7 33.4 56.7
0.75 0.0 0.1 59.9 33.4 56.5
0.75 0.2 0.1 59.1 33.5 55.7
0.75 0.5 0.1 59.7 33.6 56.5
0.0 0.1 0.1 59.4 31.9 55.8
0.25 0.1 0.1 60.1 33.2 56.6
0.5 0.1 0.1 59.4 33.3 56.5
1.0 0.1 0.1 59.3 33.5 56.6
We share some ablations that indicate how we choose hyperparameter values. The most important thing is that the losses are used, and the change that causes the most different is training with $\eta=0.0$, highlighting the importance of using 10k Words data while training.